# Universal Collection of Euclidean Invariants between Pairs of Position-Orientations

### v1

@@ 1 Introduction
There are many ways to create Euclidean equivariant neural networks, but here we focus on one specific kind: models that use scalar fields $f:\mathbb{M}_{3}\to\mathbb{R}$ on position-orientation space $\mathbb{M}_{3}:=\mathbb{R}^{3}\times S^{2}$ as their feature maps. The choice is well supported since: 1) in [5, Sec.3] it is shown theoretically that scalar fields $\mathbb{M}_{3}$ provide the same expressivity as more mathematically involved networks that use $\rho:\text{O}(3)\to\operatorname{GL}(V)$ representation fields $f:\mathbb{R}^{3}\to V$ [1, 4, 12, 13], and 2) the recent state-of-the-art results in [2] show that scalar fields on $\mathbb{M}_{3}$ can be equally expressive practically.
@@ 2 Preliminaries
The 5-dimensional smooth manifold of three-dimensional position-orientations is $\mathbb{M}_{3}=\{(x,n)\in\mathbb{R}^{3}\times\mathbb{R}^{3}\mid\|n\|=1\}$. The 5-dimensional tangent space at a point $p=(x,n)\in\mathbb{M}_{3}$ is $T_{p}\mathbb{M}_{3}=\{(\dot{x},\dot{n})\in\mathbb{R}^{3}\times\mathbb{R}^{3}\mid\dot{n}\cdot n=0\}$.
@@ 3 Universality
Let $\varphi$ be a representer for $\iota$ and let $j$ be any other invariant. Now consider $F=j\circ\varphi$. One can check that, indeed, $j=F\circ\iota$, thus showing that $\iota$ is universal.
@@ 4 Dependence
A submersion is an open map [6, Prop.4.28], meaning that it maps open sets to open sets. So, given any arbitrary open subset $U\subseteq M$ the image $\iota(U)\subseteq\mathbb{R}^{n}$ is open. There is no functional relation between all elements of an open subset in $\mathbb{R}^{n}$ (such an open set is never the graph of a function), thus $\iota$ is independent on $U$. As $U$ was arbitrary, this shows that $\iota$ is everywhere independent. ∎
@@ 5 A Universal and Independent Collection of Invariants
$v_{i}\cdot v_{j}=\left[\begin{smallmatrix}1&\iota_{4}&\iota_{1}\\ \iota_{4}&1&\iota_{2}\\ \iota_{1}&\iota_{2}&\iota_{3}\end{smallmatrix}\right]_{ij},\textrm{ with }\iota_{k}=\iota_{k}(p_{1},p_{2}),$ (5)
@@ 7 Conclusion
Acknowledgments. We gratefully acknowledge the Dutch Foundation of Science NWO for funding of VICI 2020 Exact Sciences (Duits, Geometric learning for Image Analysis, VI.C.202-031 https://www.nwo.nl/en/projects/vic202031). We thank Bekkers et al. [2] for their publicly available PONITA architecture https://github.com/ebekkers/ponita.

### v2

@@ 1 Introduction
There are many ways to create Euclidean equivariant neural networks, but here we focus on one specific kind: models that use scalar fields $f:\mathbb{M}_{3}\to\mathbb{R}$ on position-orientation space $\mathbb{M}_{3}:=\mathbb{R}^{3}\times S^{2}$ as their feature maps. The choice is well supported since: 1) in [6, Sec.3] it is shown theoretically that scalar fields $\mathbb{M}_{3}$ provide the same expressivity as more mathematically involved networks that use $\rho:\text{O}(3)\to\operatorname{GL}(V)$ representation fields $f:\mathbb{R}^{3}\to V$ [1, 5, 13, 14], and 2) the recent state-of-the-art results in [2] show that scalar fields on $\mathbb{M}_{3}$ can be equally expressive practically.
@@ 2 Preliminaries
The smooth manifold of three-dimensional position-orientations is $\mathbb{M}_{3}=\{(x,n)\in\mathbb{R}^{3}\times\mathbb{R}^{3}\mid\|n\|=1\}$. The tangent space at a point $p=(x,n)\in\mathbb{M}_{3}$ is $T_{p}\mathbb{M}_{3}=\{(\dot{x},\dot{n})\in\mathbb{R}^{3}\times\mathbb{R}^{3}\mid\dot{n}\cdot n=0\}$.
@@ 3 Universality
Let $\varphi$ be a representer for $\iota$ and let $j$ be any other invariant. Now consider $F=j\circ\varphi$. One can check that $j=F\circ\iota$, thus showing that $\iota$ is universal. ∎
Remark 1
Even with a universal collection of invariants, we cannot perfectly represent arbitrary invariants in practice, as expressivity is constrained by the chosen architecture. For instance, if our invariants are continuous and we use multi-layer perceptrons with continuous activation functions to parametrize the kernels, we can only represent continuous invariants, meaning we exclude any with discontinuities.
@@ 4 Dependence
A submersion is an open map [7, Prop.4.28], meaning that it maps open sets to open sets. So, given any arbitrary open subset $U\subseteq M$ the image $\iota(U)\subseteq\mathbb{R}^{n}$ is open. There is no functional relation between all elements of an open subset in $\mathbb{R}^{n}$ (such an open set is never the graph of a function), thus $\iota$ is independent on $U$. As $U$ was arbitrary, this shows that $\iota$ is everywhere independent. ∎
Remark 3
In practice, constructing a network using a dependent collection of invariants poses no issues, as the redundancy does not hinder training and may even enhance stability. However, proving the independence of a universal collection remains valuable, as it implies that removing any invariant renders the collection non-universal: the invariant that is removed can not be written as a function of the remaining invariants.
@@ 5 A Universal and Independent Collection of Invariants
$v_{i}\cdot v_{j}=\left[\begin{matrix}1&\iota_{4}&\iota_{1}\\ \iota_{4}&1&\iota_{2}\\ \iota_{1}&\iota_{2}&\iota_{3}\end{matrix}\right]_{ij},\textrm{ with }\iota_{k}=\iota_{k}(p_{1},p_{2}),$ (5)
@@ 6 Experiments
On some targets ($\langle R^{2}\rangle$, ZPVE, and $U_{0}$) the change in accuracy is less pronounced, in two cases even deteriorating. One explanation is that a universal collection may be unnecessary for these targets, meaning the Bekkers et al. invariants are already sufficiently expressive. The remaining variability in accuracy can be attributed to initialization and training.
On other targets ($U$, $H$, and $G$) the improvement in accuracy is stark. We can identify this as a direct consequence of using a universal collection of invariants: these targets cannot be accurately predicted using the incomplete invariants of Bekkers et al.
@@ 7 Conclusion
Acknowledgments. The Dutch Research Council (NWO) is gratefully acknowledged for financial support via VIC.202.031 (Duits, Geometric learning for Image Analysis https://www.nwo.nl/en/projects/vic202031). The European Union is gratefully acknowledged for financial support through project REMODEL (Horizon Europe, MSCA-SE, 101131557 https://doi.org/10.3030/101131557). We thank Bekkers et al. [2] for their publicly available PONITA architecture https://github.com/ebekkers/ponita.
Addendum. We thank Johan Edstedt for bringing [4, Eq.2,Prop.1] to our attention. It presents the same invariants and a similar proof of their universality. Unfortunately, we were not aware of this paper at the time of publication.